Flint Company sells many products. Gizmo is one of its popular items. Below is an analysis of the inventory purchases and sales of Gizmo for the month of March. Flint Company uses the perpetual inventory system.

Purchases Sales
Units Unit Cost Units Selling Price/Unit

3/1

Beginning inventory

130 $55

3/3

Purchase

78 $60

3/4

Sales

78 $120

3/10

Purchase

260 $65

3/16

Sales

117 $130

3/19

Sales

91 $130

3/25

Sales

65 $130

3/30

Purchase

52 $75

Using the FIFO assumption, calculate the amount charged to cost of goods sold for March.

Cost of goods sold $enter Cost of goods sold in dollars

eTextbook and Media

  

  

Using the FIFO assumption, calculate the value of ending inventory for March.

Ending Inventory $enter Ending Inventory in dollars

eTextbook and Media

  

  

Using the moving average cost method, calculate the amount assigned to the inventory on hand on March 31. (Round average cost per unit to 3 decimal places, e.g. 5.254 and final answer to 2 decimal places, e.g. 2,520.56.)

Amount assigned to the inventory enter Amount assigned to the inventory in dollars

eTextbook and Media

  

  

Using the LIFO assumption, calculate the amount assigned to the inventory on hand on March 31.

Amount assigned to the inventory $enter Amount assigned to the inventory in dollars

eTextbook and Media

  

  

Using the LIFO assumption, calculate the amount charged to cost of goods sold for March.

Cost of goods sold $enter Cost of goods sold in dollars
Answer #1
FIFO method states that goods purchased first are sold first
LIFO Method states that goods purchased later are sold first
Moving average method uses average cost for the purpose of calculation
Number of units sold = 78+117+91+65 = 351 units
Units in ending inventory = 169 units
Cost of Goods sold under FIFO =130*55+78*60+143*65 21125
Value of Ending Inventory =130*55+78*60+260*65+52*75-21125 11505
Moving Average
Cost of Goods sold
3/4 4436.25
3/16-3/25 20070.96
Total COGS 24507.21
Value of ending Inventory 8122.79
Ending Inventory as per LIFO 9685
Cost of Goods sold 22945